OpenStudy (anonymous):

Which of the following is equivalent to the statement below? If 3x + 1 = 10, then x ≠ 2.

OpenStudy (anonymous):

yes that statement is correct

OpenStudy (anonymous):

what do you mean, which of the following? But ya, the statement is right...

OpenStudy (anonymous):

A) If x ≠ 2, then 3x + 1 = 10. C)If 3x + 1 = 10, then x = 2. B)If x = 2, then 3x + 1 = 10. D)If x = 2, then 3x + 1 ≠ 10.

OpenStudy (anonymous):

ooohh, gotcha one sec

OpenStudy (anonymous):

haha alright :) take your time

OpenStudy (anonymous):

The answer would be A, because in both the equations, x does not equal 2 and the actual equation itself is 3x+1=10

OpenStudy (anonymous):

alright thank you :D

OpenStudy (anonymous):

no prob